OpenStudy (anonymous):

Mr. Shaw graphs the function f(x) = –5x + 2 for his class. The line contains the point (-2, 12). What is the point-slope form of the equation of the line he graphed? a.) y – 12 = –5(x + 2) b.) y – 12 = 2(x + 2) c.) y + 12 = 2(x – 2) d.) y + 12 = –5(x – 2)

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

y1 = y coordinate of point x1 = x coordinate of point m = slope

OpenStudy (anonymous):

where the y1 coordinate would be 12 and x1 would be -2. the slope is given in the beginning equation

OpenStudy (anonymous):

okay so would the answer be B?

OpenStudy (anonymous):

is the answer B?

OpenStudy (anonymous):

mm close.. y=mx+b where (mx) is the slope...

OpenStudy (anonymous):

the slope is -5

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

@chlobohoe @peachpi

OpenStudy (anonymous):

yep its A!

OpenStudy (anonymous):

yes